OpenStudy (anonymous):

Find a formula for the general term An of the sequence, assuming that the pattern of the first few terms continues. (5,8,11,14,17,...)

OpenStudy (kamille):

Well, you know that formula of arithmetical sequance is \[a _{n}=a _{1}+(n-1)d\] a1=5 a2=8 d can be found d=a2-a1=8-5=3 so you get \[a _{n}=a _{1}+(n-1)d=5+(n-1)3=5+3n-3=\]